# Broadband biphoton source for quantum optical coherence tomography based on a Michelson interferometer

### v2

@@ II.1 Theory
Figure 1: Experimental setup. a) Broadband biphoton source: $L$ – crystal length, $W$ – modes waist, $\Theta_{\text{p}}$ – pump walk-off angle, L1, L2 – lenses. b) QOCT Michelson interferometer: BS – beam splitter, DM – longpass dichroic mirror splitting biphotons at 1000 nm, M1 – reference mirror, M2 – sample mirror, standing on the piezo-translator which is attached to a mechanical translator, D1, D2 – single-photon detectors (SPD), based on Si avalanche photodiodes (APD) Laser components COUNT NIR, D3 – SPD, based on InGaAs APD (MPD PDM–IR).
@@
Here, $\vec{E_{j}}$ is a field amplitude, $\Theta_{j}$ is a walk-off angle, and $W_{j}$ is a beam waist. Diffraction terms are omitted under the assumption that the Rayleigh length is significantly larger than other longitudinal scales of the system. According to [38], the photon pair rate in the selected signal and idler modes can be calculated using a mode overlap integral:
$R=\frac{8d^{2}P}{\pi^{6}W^{2}_{\text{p}}W^{2}_{\text{s}}W^{2}_{\text{i}}c^{3}}\frac{\omega_{\text{s}}\omega_{\text{i}}}{n_{\text{p}}n_{\text{s}}n_{\text{i}}}\int_{-\infty}^{+\infty}\text{d}\omega\\ \times\left|\int_{-\infty}^{+\infty}\text{d}x\text{d}y\int_{-L/2}^{L/2}\text{d}zg_{\text{p}}(\vec{r})g_{\text{s}}^{*}(\vec{r})g_{\text{i}}^{*}(\vec{r})\right|^{2},$ (4)
@@
For the BBO crystal pumped at 405 nm, as studied in [37] and used in our experiment, the spectral coincidence efficiency is $S_{0}/P=125\ \frac{\text{cps}}{\text{THz}\times\text{mW}}$, and the total spectral bandwidth is given by $B_{\omega}=\frac{\kappa}{\sqrt{W}}$, where $\kappa=2\pi\times 298\ \text{THz}\times\sqrt{\mu\text{m}}$.
@@ II.2 Implementation
The alignment of the pump and target beams was meticulously optimized to achieve the maximum coincidence count rate, measured directly after the output of the fiber, reaching a final value of 3 kcps. The pump power, measured in front of the crystal, was 8 mW. Post-experiment, the crystal was removed, and the pump waist in free space was measured as $W_{p}=5.7~{}\upmu\rm m$.
@@ III.1 Theory
${{\left|f({{\omega}_{\text{s}}},{{\omega}_{\text{i}}})\right|}^{2}}=\frac{1}{\pi\delta\Delta}\underbrace{\operatorname{e}^{-\frac{{{\left({{\omega}_{1}}-{{\omega}_{2}}\right)}^{2}}}{2{{\Delta}^{2}}}}}_{\text{Phasematching}}\underbrace{\operatorname{e}^{-\frac{{{\left({{\omega}_{1}}+{{\omega}_{2}}-2{{\omega}_{0}}\right)}^{2}}}{2{{\delta}^{2}}}}}_{\text{Pump}}.$ (7)
@@ III.2 Implementation
To broaden the detection bandwidth and cover both the visible (VIS) and infrared (IR) ranges, we employed a longpass dichroic mirror (DM) with a cut-on wavelength $\lambda_{\text{c}}=1000~{}\text{nm}$. In both channels, biphotons were coupled with single-mode fibers using parabolic mirrors. The visible part was further split by a fiber beam splitter and directed to two Si-based avalanche photodiode (APD) detectors (Laser Components COUNT NIR), labeled as D1 and D2. The infrared part was directed to an InGaAs APD detector (MPD PDM-IR), labeled as D3. All detector outputs were connected to coincidence circuits, enabling coincidence measurements between the two visible-range detectors D1&D2 (VIS-VIS) and between the IR detector and both visible-range detectors: D3&D1 and D3&D2 (IR-VIS).
@@ IV Experimental results and discussion
The absolute values of their Fast-Fourier Transform (FFT) spectra are shown in Fig. 4c as thin curves with the same colors. The raw spectra qualitatively match the theory presented in Fig. 3b (see the detailed analysis in Supplementary [43]). They exhibit three peaks corresponding to the terms $\tilde{M}_{0}$, $\tilde{M}_{1}$, and $\tilde{M}_{2}$, but due to the high bandwidth, the $\tilde{M}_{0}$ and $\tilde{M}_{1}$ peaks are partially overlapped. To extract the $\tilde{M}_{0}$ term, we used the spectral range $30-\omega_{p}/$ 3, and the range $\omega_{p}/3-0.75\omega_{p}$ for extracting the $\tilde{M}_{1}$ term. To account for the spectral dependencies of the detectors’ quantum efficiencies, $\eta_{\text{VIS}}(\omega)$ and $\eta_{\text{IR}}(\omega)$, we divided the raw spectra by the corresponding efficiency products (see details in Supplementary [43]).
The processed spectrum for the $\tilde{M}_{1}$ term, combined from both VIS-VIS and IR-VIS data, is plotted as a bold dotted curve. Its Gaussian fit, shown as a gray dashed curve, has a bandwidth $\text{FWHM}_{\omega}=2\pi\times 136~{}\text{THz}$, corresponding to a wavelength range of approximately 684–992 nm. This bandwidth is close to the estimated theoretical value of $\sqrt{\frac{4\log 2}{\pi}}B_{\omega}=2\pi\times 117$ THz (5), calculated for the measured pump waist $W=5.7\ \rm\mu m$.
Considering the spectral efficiency of the detectors and the bandwidth of coincidence detection (detailed information is provided in the Supplementary [43]), the registered coincidence count rate of 3 kcps suggests a total source generation rate of $2.7\ \frac{\text{kcps}}{\text{mW}}$. The source’s spectral generation rate of $S_{0}\approx 19~{}\frac{\text{cps}}{\text{THz}\times\text{mW}}$ is about 6.6 times smaller than the theoretical value of $125~{}\frac{\text{cps}}{\text{THz}\times\text{mW}}$ [37]. The mismatch between the experimental and theoretical results can be explained by various aberrations and the wavelength dependency of the target mode waist, which were not accounted for in the theory. However, this efficiency was sufficient to achieve a registered coincidence count rate of about 25–30 $\frac{\text{cps}}{\text{mW}}$ in the raw interferogram, whereas previous works reported results of 2–4 $\frac{\text{cps}}{\text{mW}}$ for similar QOCT experiments [35, 46].
@@
Similarly, we divided the coincidence spectra related to the $\tilde{M}_{0}$ term by the products of quantum efficiencies. The result is shown in Fig. 4c as thick solid curves. Its inverse FFT, corresponding to the HOM peak with an additional sine term $M_{1}$, is shown in Fig. 4b as a blue curve. The FWHM of its Gaussian fit (gray dashed curve), which defines the QOCT resolution, was estimated to be approximately $0.46~{}\upmu\rm m$, slightly better than the record QOCT resolution of $0.54~{}\upmu\rm m$ reported in [46]. This value is consistent with the value $2\log(2)c/\text{FWHM}_{\omega}=0.49\ \upmu$m, corresponding to the measured spectral bandwidth.
@@ V Conclusion
We have introduced a cost-effective technique for generating a bright (2.7 kcps/mW pair generation rate) and broadband (136 THz bandwidth) biphoton field using a bulk nonlinear crystal with tight pump beam focusing. Demonstrating both theoretical insights and experimental validation, we applied this technique to quantum optical coherence tomography (QOCT), achieving an impressive axial resolution of $0.46\ \upmu$m, comparable to the record results reported in [32]. Significantly, we have demonstrated for the first time that Michelson interferometer-based QOCT exhibits dispersion cancellation, analogous to that observed in traditional Hong-Ou-Mandel-based implementations, resulting in at least a 50-fold improvement in resolution compared to classical Optical Coherence Tomography technique. Furthermore, our results suggest that the achieved QOCT resolution is limited by interferogram filtering procedures, indicating potential for further enhancement through the exploration of advanced filtering methods or modified interference schemes.

### v3

@@ II.1 Theory
Figure 1: Experimental setup. a) Broadband biphoton source: $L$, crystal length; $W$, modes waist; $\Theta_{\text{p}}$, pump walk-off angle; L1, L2, lenses. b) QOCT Michelson interferometer: BS, beam splitter; DM, long-pass dichroic mirror splitting biphotons at 1000 nm; M1, reference mirror; M2, sample mirror, standing on the piezotranslator, which is attached to a mechanical translator; D1, single-photon detector (SPD), based on InGaAs avalanche photodiode (APD), MPD PDM–IR; D2, D3, SPDs, based on Si APDs, Laser Components COUNT NIR.
@@
Here, $\vec{E_{j}}$ is a field amplitude, $\Theta_{j}$ is a walk-off angle, and $W_{j}$ is a beam waist. According to [38], the photon pair rate in the selected signal and idler modes can be calculated using a mode overlap integral:
$R=\frac{8d^{2}P}{\pi^{6}W^{2}_{\text{p}}W^{2}_{\text{s}}W^{2}_{\text{i}}c^{3}}\frac{\omega_{\text{s}}\omega_{\text{i}}}{n_{\text{p}}n_{\text{s}}n_{\text{i}}}\int_{-\infty}^{+\infty}\text{d}\omega_{\text{s}}\\ \times\left|\int_{-\infty}^{+\infty}\text{d}x\text{d}y\int_{-L/2}^{L/2}\text{d}zg_{\text{p}}(\vec{r})g_{\text{s}}^{*}(\vec{r})g_{\text{i}}^{*}(\vec{r})\right|^{2},$ (4)
@@
In (3), we do not account for beam divergence, assuming a Rayleigh length $z_{R}\gg L_{\text{eff}}$, which is equivalent to the inequality $W\gg\lambda_{\text{p}}/(\pi\Theta_{\text{p}})$. For the BBO crystal pumped at $\lambda_{\text{p}}=405$ nm with collinear degenerate type-I phase matching, as studied in [37] and implemented in our experiment, the walk-off angle is $\Theta_{\text{p}}=3.9^{\circ}$, and the waist should satisfy the inequality $W\gg 2~{}\upmu\rm m$.
@@
For the case described in [37] and implemented in our experiment, the spectral coincidence efficiency is $S_{0}/P=125\ \frac{\text{cps}}{\text{THz}\times\text{mW}}$, and the total spectral bandwidth is given by $B_{\omega}=\frac{\kappa}{\sqrt{W}}$, where $\kappa=2\pi\times 298\ \text{THz}\times\sqrt{\upmu\rm m}$.
@@ II.2 Implementation
The alignment of the pump and target beams was meticulously optimized to achieve the maximum coincidence count rate, measured directly after the output of the fiber, reaching a final value of 3 kcps. The pump power, measured in front of the crystal, was 8 mW. Post-experiment, the crystal was removed, and the pump waist in free space was measured as $W_{p}=5.7~{}\upmu\rm m$. This value is close to the critical threshold of $2~{}\upmu\rm m$, beyond which beam divergence should be accounted for. This could be a potential source of mismatch between the theoretical and experimental results.
@@ III.1 Theory
${{\left|f({{\omega}_{\text{s}}},{{\omega}_{\text{i}}})\right|}^{2}}=\frac{1}{\pi\delta\Delta}\underbrace{\operatorname{e}^{-\frac{{{\left(\omega_{\text{i}}-\omega_{\text{s}}\right)}^{2}}}{2{{\Delta}^{2}}}}}_{\text{Phasematching}}\underbrace{\operatorname{e}^{-\frac{{\left(\omega_{\text{i}}+\omega_{\text{s}}-2\omega_{0}\right)}^{2}}{2{{\delta}^{2}}}}}_{\text{Pump}}.$ (7)
@@ III.2 Implementation
To broaden the detection bandwidth and cover both the visible (VIS) and infrared (IR) ranges, we employed a longpass dichroic mirror (DM) with a cut-on wavelength $\lambda_{\text{c}}=1000~{}\text{nm}$. In both channels, biphotons were coupled with single-mode fibers using parabolic mirrors. The infrared part was directed to an InGaAs avalanche photodiode (APD) detector (MPD PDM-IR), labeled as D1. The visible part was further split by a fiber BS and directed to two Si-based APD detectors (Laser Components COUNT NIR), labeled as D2 and D3. All detector outputs were connected to coincidence circuits, enabling coincidence measurements between the two visible-range detectors D2&D3 (VIS-VIS) and between the IR detector and both visible-range detectors: D1&D2 and D1&D3 (IR-VIS).
@@ IV Experimental results and discussion
The absolute values of their Fast-Fourier Transform (FFT) spectra are shown in Fig. 4c as thin curves with the same colors. The raw spectra qualitatively match the theory presented in Fig. 3b (see the detailed analysis in Supplementary [43]). They exhibit three peaks corresponding to the terms $\tilde{M}_{0}$, $\tilde{M}_{1}$, and $\tilde{M}_{2}$, but due to the high bandwidth, the $\tilde{M}_{0}$ and $\tilde{M}_{1}$ peaks are partially overlapped. To extract the $\tilde{M}_{0}$ term, we used the spectral range $30-\omega_{\text{p}}/$ 3, and the range $\omega_{\text{p}}/3-3\omega_{\text{p}}/4$ for extracting the $\tilde{M}_{1}$ term. To account for the spectral dependencies of the detectors’ quantum efficiencies, $\eta_{\text{VIS}}(\omega)$ and $\eta_{\text{IR}}(\omega)$, we divided the raw spectra by the corresponding efficiency products (see details in Supplementary [43]).
The processed spectrum for the $\tilde{M}_{1}$ term, combined from both VIS-VIS and IR-VIS data, is plotted as a bold dotted curve. Its Gaussian fit, shown as a gray dashed curve, has a bandwidth $\text{FWHM}_{\omega}=2\pi\times 136~{}\text{THz}$. Taking into account possible spectral broadening caused by interferometer instability (see details in Supplementary [43]), the actual bandwidth value could be slightly lower, $\text{FWHM}_{\omega}^{\text{(corrected)}}=2\pi\times 132~{}\text{THz}$, corresponding to a wavelength range of approximately 687–986 nm. This bandwidth is close to the estimated theoretical value of $\sqrt{\frac{4\ln 2}{\pi}}B_{\omega}=2\pi\times 117$ THz (5), calculated for the measured pump waist $W=5.7\ \upmu\rm m$.
@@
Considering the spectral efficiency of the detectors and the bandwidth of coincidence detection (detailed information is provided in the Supplementary [43]), the registered coincidence count rate of 3 kcps suggests a total source generation rate of $2.7\ \frac{\text{kcps}}{\text{mW}}$. The source’s spectral generation rate of $S_{0}\approx 20~{}\frac{\text{cps}}{\text{THz}\times\text{mW}}$ is about 6 times smaller than the theoretical value of $125~{}\frac{\text{cps}}{\text{THz}\times\text{mW}}$ [37]. The mismatch between the experimental and theoretical results can be explained by various aberrations and the wavelength dependency of the target mode waist, which were not accounted for in the theory. However, this efficiency was sufficient to achieve a registered coincidence count rate of about 25–30 $\frac{\text{cps}}{\text{mW}}$ in the raw interferogram, whereas previous works reported results of 2–4 $\frac{\text{cps}}{\text{mW}}$ for similar QOCT experiments [35, 46].
Similarly, we divided the coincidence spectra related to the $\tilde{M}_{0}$ term by the products of quantum efficiencies. The result is shown in Fig. 4c as thick solid curves. Its inverse FFT, corresponding to the HOM peak with an additional sine term $M_{1}$, is shown in Fig. 4b as a blue curve. The FWHM of its Gaussian fit (gray dashed curve), which defines the QOCT resolution, was estimated to be approximately $0.46~{}\upmu\rm m$, slightly better than the record QOCT resolution of $0.54~{}\upmu\rm m$ reported in [46]. This value is consistent with the value $2\ln(2)c/\text{FWHM}_{\omega}=0.50\ \upmu$m, corresponding to the measured spectral bandwidth.
@@ V Conclusion
We have introduced a cost-effective technique for generating a bright (2.7 kcps/mW pair generation rate) and broadband (132 THz bandwidth) biphoton field using a bulk nonlinear crystal with tight pump beam focusing. Demonstrating both theoretical insights and experimental validation, we applied this technique to quantum optical coherence tomography (QOCT), achieving an impressive axial resolution of $0.46\ \upmu$m, comparable to the record results reported in [32]. Significantly, we have demonstrated for the first time that Michelson interferometer-based QOCT exhibits dispersion cancellation, analogous to that observed in traditional Hong-Ou-Mandel-based implementations, resulting in at least a 50-fold improvement in resolution compared to classical Optical Coherence Tomography technique. Furthermore, our results suggest that the achieved QOCT resolution is limited by interferogram filtering procedures, indicating potential for further enhancement through the exploration of advanced filtering methods or modified interference schemes.
Acknowledgements.
The authors would like to thank the anonymous referees for their insightful questions and constructive comments, which helped to significantly improve the quality of this work.